# BioT5+: Towards Generalized Biological Understanding with IUPAC Integration and Multi-task Tuning

### v1

@@ F.1 Molecule-oriented Cases
Input
[C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl][C]
@@
Target SMILES
Cc1noc(-c2ccc(Cl)cc2C=O)n1
Target SELFIES
[C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl]
@@
Galactica
CC1=CC=C(C2=CC(C3=CC=C(Cl)C=C3Cl)N3NN=NC3=N2)C=C1
Mol-Instructions
[C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl][C][=C][Ring1]
Text+Chem T5
CC1=NOC(C2=CC=C(Cl)C=C2C=O)=N1
BioT5+
[C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl]
@@
Input
[C][C][C][C][C][=C][C][=C][Branch1][P][C][=C][C][=C][Branch1][#Branch1][C][=Branch1]
@@
Target SMILES
BrCc1ccccc1.Oc1ccc(I)nc1Cl
Target SELFIES
[Br][C][C][=C][C][=C][C][=C][Ring1][=Branch1].[O][C]
@@
Galactica
CC1=CC=C(C2=CC=C(C(=O)O)C=C2)C=C1S(=O)(=O)[O-]
Mol-Instructions
[C][C][C][C][C][=C][C][=C][Branch1][=Branch1][C][=C][Ring1][=Branch1]
@@
Text+Chem T5
CCCCC1=CC=C(C2=CC=C(C(=O)O)C=C2)C=C1
BioT5+
[Br][C][C][=C][C][=C][C][=C][Ring1][=Branch1].[O][C]
@@
F.2 Description-guided Protein Design
@@
Input
1. The designed protein should have at least Helical transmembrane domains to ensure proper localization.
@@
Target FASTA
MSTNPKPQRKTKRNTNRRPQDVKFPGGGQIVGGVYLLPRRGPRLGVRATRKTSERSQP
@@
Input
1. The protein’s functional design necessitates the inclusion of a PxxP; stabilizes the interaction of NEF/MHC-I with host AP1 M1 ; necessary for MHC-I internalization motif.
@@
Target FASTA
MGGKWSKSSIVGWPAVRERIRQTEPRTEPAAEGVGAASQDLDKYGALTSSNTAHNNADC

### v2

@@ H.1 Molecule-oriented Cases
Input [C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl][C]
@@
Target SMILES Cc1noc(-c2ccc(Cl)cc2C=O)n1
Target SELFIES [C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl]
@@
Galactica CC1=CC=C(C2=CC(C3=CC=C(Cl)C=C3Cl)N3NN=NC3=N2)C=C1
Mol-Instructions [C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl][C][=C][Ring1]
Text+Chem T5 CC1=NOC(C2=CC=C(Cl)C=C2C=O)=N1
BioT5+ [C][C][=N][O][C][Branch1][=C][C][=C][C][=C][Branch1][C][Cl]
@@
Input [C][C][C][C][C][=C][C][=C][Branch1][P][C][=C][C][=C][Branch1][#Branch1][C][=Branch1]
@@
Target SMILES BrCc1ccccc1.Oc1ccc(I)nc1Cl
Target SELFIES [Br][C][C][=C][C][=C][C][=C][Ring1][=Branch1].[O][C]
@@
Galactica CC1=CC=C(C2=CC=C(C(=O)O)C=C2)C=C1S(=O)(=O)[O-]
Mol-Instructions [C][C][C][C][C][=C][C][=C][Branch1][=Branch1][C][=C][Ring1][=Branch1]
@@
Text+Chem T5 CCCCC1=CC=C(C2=CC=C(C(=O)O)C=C2)C=C1
BioT5+ [Br][C][C][=C][C][=C][C][=C][Ring1][=Branch1].[O][C]
@@
H.2 Description-guided Protein Design
@@
Input 1. The designed protein should have at least Helical transmembrane domains to ensure proper localization.
@@
Target FASTA MSTNPKPQRKTKRNTNRRPQDVKFPGGGQIVGGVYLLPRRGPRLGVRATRKTSERSQP
@@
Input 1. The protein’s functional design necessitates the inclusion of a PxxP; stabilizes the interaction of NEF/MHC-I with host AP1 M1 ; necessary for MHC-I internalization motif.
@@
Target FASTA MGGKWSKSSIVGWPAVRERIRQTEPRTEPAAEGVGAASQDLDKYGALTSSNTAHNNADC